# A Nonparametric Bayesian Local-Global Model for Enhanced Adverse Event Signal Detection in Spontaneous Reporting System Data

### v1

@@ 3.1 A conceptual overview on the DP Poisson model
$\lambda_{ij}\sim D_{j}\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}};\ D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j}))$ (1)
for the signal strength parameters $\{\lambda_{1j}\dots,\lambda_{Ij}\}$ associated with each drug $j$. Here $\delta_{\{\theta\}}$ denotes a degenerate distribution at $\theta$ and DP$(\alpha,G(\beta))$ is a Dirichlet process with concentration hyperparameter $\alpha>0$ and base distribution $G(\beta)\equiv$ Gamma(shape = $\beta$, rate = $\beta$) with hyperparameter $\beta>0$. The discrete mixture structure ensures that $D_{j}$, the prior for $\lambda_{ij}$, is almost surely discrete/categorical placing probability weights $\eta_{hj}$; $10\leq\eta_{hj}\leq 1;\sum_{h=1}^{\infty}\eta_{hj}=$ 1 on atoms $\theta_{hj}$. This permits clustering, allowing multiple $\lambda_{ij}$ to share an atom $\theta_{hj}$ as a common value. The probability weights $\eta_{hj}$ are generated using stick-breaking (Sethuraman,, 1994):
$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj})\text{ for }h\geq 2;\ v_{hj}\sim\text{independent Beta}(1,\alpha_{j})\text{ for }h\geq 1$ (2)
@@
based on the DP concentration hyperparameter $\alpha_{j}>0$. Meanwhile, the atoms are generated independently from the DP base distribution: $\theta_{hj}\sim G(\beta_{j})$. In practice, the infinite mixture defining $D_{j}$ can be well-approximated by a finite mixture $\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$ with a sufficiently large $K$ (Ishwaran and Zarepour,, 2002). This approximation simplifies model fitting and offers intuitive insights on the hyperparameters $\alpha_{j}$ and $\beta_{j}$, guiding their specification.
@@
Role of $\alpha_{j}$. The stick-breaking representation (2) elucidates that when $\alpha_{j}<1$, each $v_{hj}$ is, on average, greater than $1/2$ with smaller $\alpha_{j}$ favoring $v_{hj}$ closer to one. Consequently, $\eta_{hj}$ decreases sharply in $h$, leading the mixture (1), or its finite approximation with $K$ components, to retain only a few non-empty components with sufficiently large weights $\eta_{hj}$. Under a finite mixture approximation with $K_{+}<K$ non-empty components, (1) models $\lambda_{ij}$ using a categorical distribution with $K_{+}$ categories, placing point masses at the atoms $\theta_{1j},\dots,\theta_{K_{+}j}$ with weights $\eta_{1j},\dots,\eta_{K_{+}j}$. This fosters clustering, restricting $\lambda_{1j},\dots,\lambda_{Ij}$ to at most $K_{+}$ distinct values. When $K_{+}<I$, multiple $\lambda_{ij}$ share the same cluster $h$ admitting the common value $\theta_{hj}$, facilitating their pooling for information sharing. While a smaller $\alpha$ encourages a smaller $K_{+}$ aiding greater information sharing, a larger $K_{+}$ ($>2$) is permitted in the posterior when supported by data. This contrasts with the GPS framework of DuMouchel, (1999) commonly used for SRS signal detection (see Section S.1 for a review) which strictly categorizes $\lambda_{ij}$ into two groups, intuitively, signal ($\lambda_{ij}\leq 1$) and non-signal ($\lambda_{ij}\gg 1$). The DP framework, by allowing $K_{+}>2$, enhances heterogeneity acknowledgment within both signal and non-signal groups. Figure 2a illustrates this with $I=6$ AEs and $K_{+}=3$ clusters.
@@
Because a smaller $\alpha_{j}$ promotes greater information sharing, a prior assigning sufficient mass to small positive values is recommended for stable estimation of $\lambda_{ij}$. Hu et al., (2015) used a Uniform(0.2, 10) prior for $\alpha_{j}$. We suggest a more aggressive weakly informative default prior, viz., Exponential(rate = $\psi_{\alpha}$) with a moderately large $\psi_{\alpha}$, e.g., $\psi_{\alpha}=3$ (placing $\approx 95\%$ mass in $(0,1)$), which favors small $\alpha_{j}$ while allowing larger values if needed. This choice also ensures conjugacy for efficient Gibbs sampling updates for MCMC implementation.
@@
Role of $\beta_{j}$. The parameter $\beta_{j}$ controls variability among the atoms $\theta_{hj}$ drawn from the base $G(\beta_{j})\equiv$ Gamma(shape = $\beta_{j}$, rate = $\beta_{j}$) distribution, thereby controlling the shrinkage in $\lambda_{ij}$. Since $G(\beta_{j})$ has unit mean and variance $1/\beta_{j}$, larger (smaller) values of $\beta_{j}$ encourage all $\theta_{hj}$ to concentrate more (less) around $1$, inducing more (less) shrinkage in $\lambda_{ij}$ (around $1$).
@@
Remark.
[…]
@@ 3.2 The local-global DP Poisson model
The DP model in Section 3.1 enables flexible and effective signal detection for each drug separately but does not account for associations between drugs in their signal strenghts. To capture these associations, we introduce the concept of local and global (prior) distributions for AE signal strengths. Specifically, we assume that the realized AE signal strengths $\{\lambda_{1j},\dots,\lambda_{Ij}\}$ for each drug $j$ is drawn from a mixture of a “local” distribution specific to the drug and a “global” distribution shared among drugs with similar compositions or functions.
@@
$D_{j}\sim DP(\alpha_{j},G(\beta_{j}));\ D\sim DP(\alpha,G(\beta)).$ (3)
@@
We note that the proposed local-global mixture DP prior (3) enables information sharing in signal strengths across both AEs and drugs. Specifically, for each drug $j$, the realized signal strengths $\{\lambda_{i^{\prime}j}:\text{all }i^{\prime}\text{ with }z_{i^{\prime}j}=1\}$ exhibit the same $D_{j}$-induced clustering (see Section 3.1) as the local strengths $\{\lambda_{1j}^{l},\dots,\lambda_{Ij}^{l}\}$, facilitating information sharing across AEs. Meanwhile, for any AE $i$, the signal strengths $\{\lambda_{ij^{\prime}}:\text{all }j^{\prime}\text{ with }z_{ij^{\prime}}=0\}$ form a cluster attaining the common global value $\lambda_{i}^{g}$, and enable information sharing across drugs (the global values $\{\lambda_{1}^{g},\dots,\lambda_{I}^{g}\}$ themselves admitting clustering via $D$). Figure 2B illustrates this mechanism using $I=6$ AEs and $J=3$ drugs. This bi-level clustering contrasts with the single-level (AE-only) clustering induced by the DP prior of Hu et al., (2015). As demonstrated in extensive simulations (Section 4), the bi-level clustering and information sharing play a critical role in stabilizing the posterior and enhancing signal detection for $\lambda_{ij}$, particularly for drug-AE pairs where $n_{i\bullet}$ is small (rare AEs) and/or $n_{\bullet j}$ is small (newer drugs), leading to small $E_{ij}$.
The drugs under consideration should be similar in composition or function to justify a single global distribution $D$. If distinct drug groups exist with no similarities, separate global processes can be used to enable separate within-group information sharing. Additionally, SRS data often include a collapsed, heterogeneous comparator category $J$ (e.g., “Other drugs”), serving as a baseline group with few, if any, distinct AE signals. In such cases, shrinking all $\lambda_{iJ}$ toward $1$ using a prior such as $\text{Gamma}(\text{shape}=\tau,\text{rate}=\tau)$ with an appropriate hyperparameter $\tau$ can improve signal detection and inference (see Supplement S.2.1).
As noted in Remark Remark, a separate ZIP likelihood $n_{ij}\sim\text{ZIP}(\lambda_{ij}E_{ij},\omega_{j})$ is unnecessary for signal detection due to the flexibility of the DP prior. However, it may be used if direct inference on structural zero positions/probabilities is required. Complete hierarchical Bayesian representations of our Poisson and ZIP-based models are provided in Supplement S.2.1.
@@
The Hu et al., (2015) model can be viewed as a “local-only” special case of our local-global model, with only drug-specific local distributions $D_{j}$ for $\lambda_{ij}$ and no global distribution (i.e., $\pi^{l}=1$). At the other extreme, a “global-only” model ($\pi^{l}=0$) yields common $\lambda_{i}$ across drugs. This is a restrictive model that cannot differentiate between drugs and only permits “global” signal detection—assessing whether an AE-$i$ is a signal for any drug under consideration (see Supplement S.2.2 for further discussion on this model).
@@
Extensive literature exists on local-global DP models for sharing information across groups while preserving DP-based clustering within each group. A common approach employs hierarchical DPs (Teh et al.,, 2004), where local DPs share a global DP as a base distribution. In contrast, our approach leverages a two-component DP mixture framework (Antoniak,, 1974), separately modeling local and global AE signal strengths and thus offering a more transparent bi-level clustering over AEs and drugs, enabling more flexible estimation of the local and global parameters for greater flexibility in SRS applications.
@@
In this paper, we focus on drug-based pharmacovigilance, identifying AE (row) signals given one or more drugs (columns in SRS contingency tables). An alternative AE-based approach discovers drugs strongly associated with a given set of AEs (Ding et al.,, 2020; Liu et al.,, 2024). A row-based version of our column-based prior (3), obtained by switching the roles of $i$ and $j$, can be used for such analyses.

### v2

@@ 3.2 The Gamma Poisson Shrinker (GPS) model
$\lambda_{ij}\sim D_{j}\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}};\ D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j})).$ (1)
[…]
$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj})\text{ for }h\geq 2;\ v_{hj}\sim\text{independent Beta}(1,\alpha_{j})\text{ for }h\geq 1.$ (2)
@@
Meanwhile, the atoms $\theta_{hj}$ are generated independently from the DP base gamma distribution: $\theta_{hj}\sim G(\beta_{j})$; $h\geq 1$. In practice, the infinite mixture (1) defining $D_{j}$ can be well-approximated ishwaran2002exact by a finite mixture $\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$ with a sufficiently large number of components $K$; this finite mixture approximation provides a computationally tractable approach to implementing this model.
Hu et al.hu2015signal consider full Bayesian inference on $\lambda_{ij}$ by first eliciting independent vague prior distributions on the hyperparameters $\alpha_{j}$ (specifically a Uniform(0.2, 10) prior) and $\beta_{j}$ (a Uniform(0, 1) prior) and then utilizing MCMC sampling from the resulting joint posterior. Subsequently, the drug-$j$ and AE-$i$ pair is identified as a signal if the 5th posterior percentile of $\lambda_{ij}$ computed using its posterior MCMC draws exceeds a certain threshold. The authors suggest using the deterministic threshold of $2$ for this purpose; a more data-driven threshold controlling FDR at a prespecified level (e.g., $0.05$) may be obtained.ahmed2010false
Our proposed approach is motivated by this method, and we revisit this approach elucidating the key roles played by the model hyperparameters $\alpha_{j}$ and $\beta_{j}$, and guiding their prior specification for full Bayesian inference in Section 4.1.
3.4 ZIP model based Pseudo LRT (frequentist) method
In parallel to Bayesian developments, major advances on SRS data mining for AE signal detection have also been made using frequentist methods, notably using the machinery of the likelihood ratio test huang2011likelihood ; huang2013likelihood ; huang2014likelihood . These approaches often have the advantage of built-in FDR control while ensuring high sensitivity. We consider the ZIP model for the reporting count $n_{ij}\sim\text{ZIP}(n=n_{ij}\mid\lambda_{ij}\times E_{ij},\omega_{j})$. The pseudo LRT, as proposed by Chakraborty et al.chakraborty2022use , is a recent method designed to accommodate the zero-inflated nature of SRS data while easing computational demands. Similar to the standard LRT method, this approach first tests a global alternative hypothesis $H_{aj}:\lambda_{ij}>1$ for at least one $i\in\{1,\dots,I\}$ against the null hypothesis $H_{0j}:\lambda_{ij}=1$ for all $i\in\{1,\dots,I\}$ for a given drug-$j$. If $H_{aj}$ is rejected, a step-down procedure is used to conduct individual tests on $H_{aij}:\lambda_{ij}>1$.
The ordinary LRT estimates the nuisance parameter $\omega_{j}$ separately for $H_{0j}$ and $H_{0j}\cup H_{aj}$, while the pseudo LRT is constructed using a global point estimate $\hat{\omega}_{j}$ of $\omega_{j}$ consistent under $H_{0j}\cup H_{aj}$ in both the numerator and the denominator of the likelihood ratio test statistic. Chakraborty et al consider the point estimate of $\omega_{j}$ is obtained by maximizing the profile likelihood $\ell_{j}(\omega_{j})$ for this purpose, with $\hat{\lambda}_{ij}=\max\{1,n/E_{ij}\}$ as the maximum likelihood estimate of $\lambda_{ij}$. The resulting statistic is given by
$\text{LR}_{ij}=\exp\left[-(\hat{\lambda}_{ij}-1)E_{ij}\right]\hat{\lambda}_{ij}^{n_{ij}}.$
The maximum likelihood ratio statistic for drug-$j$ is then defined as $\text{MLR}_{j}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\}\end{subarray}}\text{LR}_{ij}$. In the global test, a parametric bootstrap is used to approximate the p-value $P_{j}=\Pr(\text{MLR}_{j}\geq\text{MLR}_{j}^{\text{obs}})$ using (parametric bootstrap) draws for $\text{LR}_{ij}$ under $H_{0j}$ and hence $\text{MLR}_{j}$, where $\text{MLR}_{j}^{\text{obs}}$ is the observed test statistic. If the global test is rejected, step-down p-values chakraborty2022use ; huang2011likelihood are computed for the individual tests using the same parametric bootstrap samples for $\text{MLR}_{j}$ and using the individual observed likelihood ratio test statistic values, such that $P_{ij}=\Pr(\text{LR}_{ij}\geq\text{MLR}_{ij}^{\text{obs}})$. For multiple drugs, the same approach can be applied, defining an overall maximum likelihood ratio as $\text{MLR}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\},j\in\{1,\dots,J\}\end{subarray}}\text{LR}_{ij}$. This step-down procedure employing the same global test statistic strictly controls the false discovery rate at the same level $\alpha$ of the global test performed using the MLR huang2011likelihood ; chakraborty2022use . When multiple drugs $j$ are tested simultaneously, one considers the extended version with global maximum likelihood ratio $\max_{j\in\{1,\dots,J\}}\text{MLR}_{j}$.
4 A NOVEL LOCAL-GLOBAL DP MODEL FOR SRS DATA MINING
We begin this section by revisiting the DP Poisson model of Hu et al. hu2015signal (Section 4.1), highlighting the role of its hyperparameters in SRS signal discovery with the aid of a schematic plot (Figure 2a) and providing guidance on prior specification for these hyperparameters. We then introduce our novel local-global DP framework (Section 4.2), which accommodates between-drug associations and enables effective information sharing across drugs, thereby enhancing SRS data mining.
4.1 The DP Poisson model revisited: the roles of the hyperparameters in SRS signal discovery and the effect of zero-inflation
Role and prior specification for the hyperparameter $\alpha_{j}$
@@ 4 A NOVEL LOCAL-GLOBAL DP MODEL FOR SRS DATA MINING
The stick-breaking representation (2) elucidates that when $\alpha_{j}<1$, each $v_{hj}$, on average, is greater than $1/2$ with smaller values of $\alpha_{j}$ favoring $v_{hj}$ closer to one. Consequently, $\eta_{hj}$ decreases sharply in $h$, leading the mixture (1), or its finite approximation with $K$ components, to retain only a few non-empty components with sufficiently large and positive weights $\eta_{hj}$. Under a finite mixture approximation with $K_{+}<K$ non-empty components, (1) models $\lambda_{1j},\dots,\lambda_{Ij}$ using a categorical distribution with $K_{+}$ categories, placing point masses at the atoms $\theta_{1j},\dots,\theta_{K_{+}j}$ with weights $\eta_{1j},\dots,\eta_{K_{+}j}$. This fosters clustering, restricting $\lambda_{1j},\dots,\lambda_{Ij}$ to at most $K_{+}$ distinct values. When $K_{+}<I$, multiple $\lambda_{ij}$ share the same cluster $h$ admitting the common value $\theta_{hj}$, facilitating their pooling for information sharing. It follows that a smaller $\alpha$ encourages a smaller $K_{+}$, resulting in fewer clusters and thus greater information sharing among the $\lambda_{ij}$, while a larger $\alpha$ induces a larger $K_{+}$, resulting in more clusters that collectively capture more heterogeneity among the $\lambda_{i}j$ values. Importantly, when supported by the data and induced through a (moderately) large $\alpha$, values of $K_{+}>2$ are permitted in the posterior of this method. This contrasts with the GPS method of dumouchel1999bayesian (see Section 3 for a review) which strictly categorizes $\lambda_{ij}$ into two groups, intuitively, signal ($\lambda_{ij}\leq 1$) and non-signal ($\lambda_{ij}\gg 1$). The DP framework, by allowing $K_{+}>2$, enhances heterogeneity acknowledgment within both signal and non-signal groups. Figure 2a illustrates this with $I=6$ AEs and $K_{+}=3$ clusters.
@@
Because a smaller $\alpha_{j}$ promotes greater information sharing, a prior assigning sufficient mass to small positive values is recommended for stable estimation of $\lambda_{ij}$. Hu et al.hu2015signal used a Uniform(0.2, 10) prior for $\alpha_{j}$. We suggest a more aggressive weakly informative default prior, viz., Exponential(rate = $\psi_{\alpha}$) with a moderately large $\psi_{\alpha}$, e.g., $\psi_{\alpha}=3$ (placing $\approx 95\%$ mass in $(0,1)$). This prior favors small $\alpha_{j}$ while allowing larger values if needed. This choice also ensures conjugacy for efficient Gibbs sampling updates for MCMC implementation.
@@
Role and prior specification for the hyperparameter $\beta_{j}$
The parameter $\beta_{j}$ controls variability among the atoms $\theta_{hj}$ drawn from the base $G(\beta_{j})\equiv$ Gamma(shape = $\beta_{j}$, rate = $\beta_{j}$) distribution, thereby controlling the shrinkage in $\lambda_{ij}$. Since $G(\beta_{j})$ has unit mean and variance $1/\beta_{j}$, larger (smaller) values of $\beta_{j}$ encourage all $\theta_{hj}$ to concentrate more (less) around $1$, inducing more (less) shrinkage in $\lambda_{ij}$ (around $1$).
@@
Effect of zero-inflation
Zero inflation can be handled to some extent in the DP Poisson model through a small atom $\theta_{hj}\approx 0$, causing $\lambda_{ij}$ and hence Poisson($\lambda_{ij}E_{ij}$) to concentrate near zero for all AEs $i$ in cluster $h$, effectively mimicking a degenerate-at-zero $\delta_{\{0\}}$ distribution. This yields a model similar to $n_{ij}\sim$ ZIP($\lambda_{ij}E_{ij},\omega_{j})$, $\omega_{j}$ being the structural zero probability, with an analogous DP prior on $\lambda_{ij}$, producing similar posterior for $\lambda_{ij}$. Consequently, signal detection and inference on $\lambda_{ij}$ are also expected to be broadly similar in both models (Section 5 provides numerical validation). However, some (small) differences may arise in practice, particularly when analyzing large real-world SRS datasets with complex signal patterns, due to numerical rounding errors for handling zero atoms. Furthermore, for direct inference on the structural zero positions, the Poisson approximation of $\delta_{\{0\}}$ is insufficient, and a ZIP model $n_{ij}\sim\text{ZIP}(\lambda_{ij}E_{ij},\omega_{j})$ is needed. The same prior (3) may still be used on $\lambda_{ij}$ assuming $\lambda_{ij}=0$ whenever the underlying cell-specific zero-inflation indicator, say $y_{ij}$, indicates a structural zero to ensure identifiability. For any drug-$j$, the posterior distributions of $y_{ij}$ and/or the column-specific zero-inflation probability $\omega_{j}$ provide inference on whether AE-$i$ and/or any AE is a structural zero, respectively. In practice, it is thus important to assess whether structural zeros/zero inflation are likely in the SRS dataset being analyzed, and whether inference on them is of interest, in order to guide appropriate model selection. The DP-Poisson framework offers some robustness against model misspecification due to zero inflation, but explicit ZIP modeling may be preferred when structural zeros are a primary concern.
@@ 4.2 The local-global DP Poisson model
The DP model in Section 4.1 enables flexible and effective signal detection for each drug separately but does not account for associations between drugs in their signal strengths. To capture these associations, we introduce the concept of local and global (prior) distributions for AE signal strengths. Specifically, we assume that the realized AE signal strengths $\{\lambda_{1j},\dots,\lambda_{Ij}\}$ for each drug $j$ is drawn from a mixture of a “local” distribution specific to the drug and a “global” distribution shared among drugs with similar compositions or functions.
@@
$D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j}));\ D\sim\text{DP}(\alpha,G(\beta)).$ (3)
@@
We note that the proposed local-global mixture DP prior (3) enables information sharing in signal strengths across both AEs and drugs. Specifically, for each drug $j$, the realized signal strengths $\{\lambda_{i^{\prime}j}:\text{all }i^{\prime}\text{ with }z_{i^{\prime}j}=1\}$ exhibit the same $D_{j}$-induced clustering (see Section 4.1) as the local strengths $\{\lambda_{1j}^{l},\dots,\lambda_{Ij}^{l}\}$, facilitating information sharing across AEs. Meanwhile, for any AE $i$, the signal strengths $\{\lambda_{ij^{\prime}}:\text{all }j^{\prime}\text{ with }z_{ij^{\prime}}=0\}$ form a cluster attaining the common global value $\lambda_{i}^{g}$, and enable information sharing across drugs (the global values $\{\lambda_{1}^{g},\dots,\lambda_{I}^{g}\}$ themselves admitting clustering via $D$). Figure 2B illustrates this mechanism using $I=6$ AEs and $J=3$ drugs. This bi-level clustering contrasts with the single-level (AE-only) clustering induced by the DP prior of Hu et al.hu2015signal . As demonstrated in extensive simulations (Section 5), the bi-level clustering and information sharing play a critical role in stabilizing the posterior and enhancing signal detection for $\lambda_{ij}$, particularly for drug-AE pairs where $n_{i\bullet}$ is small (rare AEs) and/or $n_{\bullet j}$ is small (newer drugs), leading to small $E_{ij}$.
The drugs under consideration should be similar in composition or function to justify a single global distribution $D$. If distinct drug groups exist with no similarities, separate global processes can be used to enable separate within-group information sharing. Additionally, SRS data often include a collapsed, heterogeneous comparator category $J$ (e.g., “Other drugs”), serving as a baseline group with few, if any, distinct AE signals. In such cases, shrinking all $\lambda_{iJ}$ toward $1$ using a prior such as $\text{Gamma}(\text{shape}=\tau,\text{rate}=\tau)$ with an appropriate hyperparameter $\tau$ can improve signal detection and inference (see Supplement S.1.1).
As noted in Section 4.1, while the DP-Poisson framework is expected to provide some robustness against zero-inflation, a separate ZIP likelihood $n_{ij}\sim\text{ZIP}(\lambda_{ij}E_{ij},\omega_{j})$ may be used if substantial zero inflation is present in the data, or direct inference on structural zero positions/probabilities is required. Complete hierarchical Bayesian representations of our Poisson and ZIP-based models are provided in Supplement S.1.1.
@@
The Hu et al.hu2015signal model can be viewed as a “local-only” special case of our local-global model, with only drug-specific local distributions $D_{j}$ for $\lambda_{ij}$ and no global distribution (i.e., $\pi^{l}=1$). At the other extreme, a “global-only” model ($\pi^{l}=0$) yields common $\lambda_{i}$ across drugs. This is a restrictive model that cannot differentiate between drugs and only permits “global” signal detection—assessing whether an AE-$i$ is a signal for any drug under consideration (see Supplement S.1.2 for further discussion on this model).
@@
Extensive literature exists on local-global DP models for sharing information across groups while preserving DP-based clustering within each group. A common approach employs hierarchical DPs teh2004sharing , where local DPs share a global DP as a base distribution. In contrast, our approach leverages a two-component DP mixture framework antoniak1974mixtures , separately modeling local and global AE signal strengths and thus offering a more transparent bi-level clustering over AEs and drugs, enabling more flexible estimation of the local and global parameters for greater flexibility in SRS applications.
@@
In this paper, we focus on drug-based pharmacovigilance, identifying AE (row) signals given one or more drugs (columns in SRS contingency tables). An alternative AE-based approach can be performed to discover drugs strongly associated with a given set of AEs ding2020evaluation ; liu2024pattern . Our proposed approach can be readily adapted for such analyses: a row-based version of our column-based prior (3), obtained by switching the roles of $i$ and $j$, can be used for this purpose.